OpenStudy (ps22271):

When radioactive substances​ decay, the amount remaining will form a geometric sequence when measured over constant intervals of time. The table shows the amount of a radioactive isotope initially and after 2 hours. What are the amounts left after 1​ hour, 3​ hours, and 4​ hours? Hours elapsed 0 1 2 3 4 Grams of the isotope 1926 501 How much of the isotope is left after 1 hour?

OpenStudy (danjs):

did you get anywhere yet?

OpenStudy (ps22271):

No, I have 3 chances to get it right and I have gotten it wrong twice already. I set up an equation and everything but I'm getting no where with it.

OpenStudy (danjs):

Recall, for geometric sequence, each next term is a constant multiple 'r' times the last term. So the general form is... \[\large y=A*r^t\] where A is the initial value, when t=0

OpenStudy (danjs):

they give you two points, (t,y) (0 , 1926) (2 , 501)

OpenStudy (danjs):

you can use those to figure out what the values of A and r will be... \[y=A*r^t\] \[\large 1926=A*r^0=A\] With that, A=1926, now use the second point.. \[y=1926*r^t\] \[\large 501 = 1926*r^2\]

OpenStudy (ps22271):

then you would get 0.2601 right?

OpenStudy (danjs):

the square root of that, so r=0.51, and the function for grams left y in terms of hours passed t, is for t=0,1,2,3.... \[\huge y(t)=1926*(0.51)^t\]

OpenStudy (danjs):

you get it all?

OpenStudy (ps22271):

so then the answer for 1 gram would be 982.26?

OpenStudy (danjs):

yeah, 1 hour elapsed, grams left 982.26

OpenStudy (ps22271):

thank you so so much

OpenStudy (danjs):

welcome

OpenStudy (ps22271):

I have one more math question I need help with. You are offered a job that pays ​$39,000 during the first​ year, with an annual increase of 10​% per year beginning in the second year. That​ is, beginning in year​ 2, your salary will be 1.1 times what it was in the previous year. What can you expect to earn in your fourth year on the​ job? Round your answer to the nearest dollar. In your fourth year on the​ job, you can expect to earn ​$_

OpenStudy (danjs):

similar to the first question...

OpenStudy (danjs):

If you want to make the first year n=1, then you need to alter it a litttle bit, to make n=1 the initial value A \[\large y=A*r^{t-1}\]

OpenStudy (danjs):

you have A=39000, and r=1.1 \[\huge y(t)=39000*(1.1)^{t-1}\]

OpenStudy (danjs):

you get that?

OpenStudy (ps22271):

Yes i do. That makes so much more sense. So then to do year 4 i would put 4 in for t and then solve and get the answer for the earning of the fourth year? you don't do anything with the 10%?

OpenStudy (danjs):

the r=1.1 value adds 10% to each previous year

OpenStudy (ps22271):

Ahhhh okay.

OpenStudy (ps22271):

So for the 4th year you'd pay $57,098.9? correct?

OpenStudy (danjs):

no, remember we had to change the exponent to make the first term A=39000, the start value after t=1. so the exponent is t-1

OpenStudy (ps22271):

so then the equation wouldn't be 39000(1.1)4-1 ?

OpenStudy (ps22271):

because wouldn't the t be replaced by 4 since it would be the 4th year?

OpenStudy (ps22271):

are you still there?

OpenStudy (danjs):

yeah sorry...

OpenStudy (danjs):

Yes, for the 4th year, you use t=4, which gives \[\large y(4)=39000*(1.1)^{4-1}\] the exponent becomes 3, you still used 1.1^4 in your calculation should be \[\large y(4)=39000*(1.1)^{4-1}=51909\]